a man is 4 times as old as his son.20 years later he will be just twice as old as his son.find their present ages.
Answer by Maths68(1474)

Let
Present age of Man = m
Present age of Son = s

Man is 4 times as old as his son.
m=4s.............(1)
20 years later
Man age = m+20
Son age = s+20
Then he will be just twice as old as his son.
m+20=2(s+20)
m+20=2s+40
m=2s+40-20
m=2s+20..............(2)
Put the value of m from (1) to (2)
4s=2s+20
4s-2s=20
2s=20
2s/2=20/2
s=10
Put the value of s in (1)
m=4s
m=4(10)
m=40

Present age of Man = m = 40 years old
Present age of Son = s = 10 years old